Financial ratios computed for Whittaker Inc. include the following:

Current ratio 1.9 to 1
Acid-test ratio 1.4 to 1
Debt/equity ratio 2.0 to 1
Inventory turnover 3.6 times
Accounts receivable turnover 5.4 times
Times interest earned 4.60 times
Gross profit ratio 40 %
Return on investment 7.17 %
Earnings per share $ 3.40
  • All sales during the year were made on account. Cash collections during the year exceeded sales by $13,000, and no uncollectible accounts were written off.
  • The balance of the accounts receivable account was $55,000 on January 1, 2020.
  • No common stock was issued during the year.
  • Dividends declared and paid during the year were $2,228.
  • The balance of the inventory account was $37,300 on January 1, 2020.
  • Interest expense on the income statement relates to the 10% bonds payable; $12,000 of these bonds were issued on May 1, 2020; the remaining amount of bonds payable were outstanding throughout the year. All bonds were issued at face amount.a. Complete the income statement and balance sheet for Whittaker Inc. Answer is not complete. WHITTAKER INC. Income Statement
  • Answer is not complete. WHITTAKER INC. Balance Sheet December 31, 2020 Current assets: Cash $ 190,000 63.000 Accounts receiva
Answer #1
WHITTAKER INC.
Income Statement
For the Year Ended December 31, 2020.
Particulars Amount
Sales (Exp. 5) $261,900
Cost of goods sold (Exp. 6) $157,140
Gross Profit (Bal. Fig.) $104,760
Operating Expenses(Bal. Fig.) $77,610
Income from Operations (Exp. 7) $27,150
Interest Expense (Exp. 8) $5,900
Income before Taxes $21,250
Income Tax (20%) $4,250
Net Income $17,000
WHITTAKER INC.
Balance Sheet
December 31, 2020.
Particulars Amount
Current Assets
      Cash (Bal. Fig.) $98,000
      Accounts Receivable, Net (Exp. 3) $42,000
      Inventory (Exp. 2) $50,000
           Total Current Assets $190,000
Property, Plant & Equipement, Net (Bal. Fig.) $54,500
                     Total Assets $244,500
Current Liabilities (Exp. 1) $100,000
Bonds Payable, 10% $63,000
           Total Liabilities $163,000
Stockholder's Equity:
      Common Stock, $4 per value $20,000
      Additional Paid-in Capital $30,000
      Retained Earnings (Exp. 4) $31,500
           Total Stockholder's Equity $81,500
                  Total Liabilities & Stockholder's Equity $244,500
Exp. 1 Current Liabilities
CA/CL = 1.9 (Current Ratio)
190000/CL = 1.9
CL = 190000/1.9
Current Liabilities = $ 100000
Exp. 2 Inventory
CA-Inventory/CL = 1.4 (Acid Test Ratio)
190000-Inv/100000 = 1.4
Inventory = 190000-140000
Inventory = $ 50000
Exp. 3 Accounts Receivable
Since all the sales have been made on the account, therefore there is no increase in the receivable accounts. While the Cash received exceeds the total sales by $ 13000, that means the $ 1300 is received from opening receivables.
Accounts Receivable = Opening Receivables - 13000
Accounts Receivable = 55000-13000
Accounts Receivable = $ 42000
Exp. 4 Retained Earnings
Debt/Equity = 2 (Debt Equity Ratio)
163000/Equity = 2
Equity = 81500
Retained Earnings = Equity - Capital
Retained Earnings = 81500 - (20000+30000)
Retained Earnings = $ 31500
Exp. 5 Sales
Sales/Average Receivables = 5.4 (Accounts Receivable Turnover Ratio)
Sales/(55000+43000/2) = 5.4
Sales/48500 = 5.4
Sales = $ 261900
Exp. 6 Cost of Goods Sold
Cost of Goods Sold/ Average Inventory = 3.6 (Inventory Turnover Ratio)
Cost of Goods Sold/ (37300+50000/2) = 3.6
Cost of Goods Sold/ 43650 = 3.6
Cost of Goods Sold = $ 157140
Exp. 7 Interest Expense
= ((63000-12000)*10%) + (12000*10%*8/12)
= (51000*10%) + (12000*10%*8/12)
= 5100 + 800
Interest Expense = $ 5900
Exp. 8 Income from Operations
Income from Operations(EBIT)/Interest Expense = 4.6 (Times Interest Earned)
Income from Operations(EBIT)/5900 = 4.6
Income from Operations(EBIT) = $ 27150